# An information-matching approach to optimal experimental design and active learning

### v2

@@ Abstract
Many OED criteria utilize the Fisher information matrix (FIM), whose inverse establishes a lower bound on parameter covariance, known as the Cramér-Rao bound [22, 23, 24]. Common approaches optimize parameter precision through the FIM, such as by minimizing its trace (A-optimality) [25, 26, 27], or maximizing its determinant (D-optimality) [28, 29, 16, 27] or smallest eigenvalues (E-optimality) [30, 29, 31]. However, many applications of predictive models do not require precise parameter estimates per se, but accurate predictions for key quantities of interest (QoIs) [18, 32, 33, 34, 35]. This distinction is well illustrated by sloppy models, where many parameter combinations are practically unidentifiable, yet still yield precise predictions [36, 37, 38]. Sloppy models exhibit a characteristic information spectrum, where FIM eigenvalues are nearly uniformly spaced on a log scale over many orders of magnitude. Many eigenvalues are smaller than the model’s evaluation precision, rendering the above OED criteria sensitive to numerical noise.
Motivated by these considerations, we develop an information-matching method for OED that prioritizes the precision of predictions for the target QoIs. This approach uses a model parameterized by $\boldsymbol{\theta}$ in two key scenarios: training, $\bm{f}$, and prediction, $\bm{g}$ (see Fig. 1). Given a dataset of $M$ independent inputs $\{\mathbf{x}_{m}\}_{m=1}^{M}$ and their corresponding ground truth labels $\{\mathbf{p}_{m}\}_{m=1}^{M}$, we first use $\bm{f}(\boldsymbol{\theta};\mathbf{x}_{m})$ to train the parameters against the data $\{\mathbf{x}_{m},\mathbf{p}_{m}\}$. Then, we use the trained parameters to predict the target QoIs $\mathbf{q}$ corresponding to the input $\mathbf{y}$ through the mapping $\bm{g}(\boldsymbol{\theta};\mathbf{y})$. In this scenario, $\mathbf{y}$ may act as an input, similar to $\mathbf{x}_{m}$, as a control parameter for the QoIs, or as a discrete index to distinguish between different QoI values. The information-matching method leverages both scenarios to identify the minimal subset of training data that contains the information necessary to precisely constrain the parameters relevant to target QoIs. Our strategy is, given a target precision for target QoIs, align the FIM for the training data with that of the QoIs, ensuring the training data carry the information needed to constrain the predictions precisely. Thus, only the parameter combinations that need to be identified are trained, bypassing numerical stability issues in cases where the FIMs are ill-conditioned.
@@
where the weight $w_{m}$ is the inverse variance of the label $\mathbf{p}_{m}$, i.e., $w_{m}=1/\sigma_{m}^{2}$. As shown in the supplement, the FIM for the training scenario is given by
@@
Let $\bm{g}(\boldsymbol{\theta};\mathbf{y})$ denote a mapping from the model parameters $\boldsymbol{\theta}$ to the target QoIs for input $\mathbf{y}$ that is analytic at $\boldsymbol{\theta}_{0}=\left<\boldsymbol{\theta}\right>_{\boldsymbol{\theta}}$, where $_{\boldsymbol{\theta}}$ denotes an expectation value over the distribution of parameters. Consider parameters of the form $\boldsymbol{\theta}=\boldsymbol{\theta}_{0}+\epsilon\delta\boldsymbol{\theta}$. If the constraints in Eq. (4) are satisfied, then
@@
Theorem 1 states that the uncertainties of the QoIs propagated from the optimal training data [$\mathrm{Cov}(\bm{g})$] are within the predefined target uncertainties ($\mathbf{\Sigma}$), up to third order in $\epsilon$. The information-matching method is unique in that it simultaneously aims to minimize data usage while ensuring adequate information for precise predictions. In contrast, previously mentioned OED criteria (A-, D-, and E-optimality) only prioritize reducing some measure of parameter variance. By minimizing the number of data points, the information-matching approach not only enhances efficiency but also improves model interpretability by focusing the analysis on only the most critical training data.
@@
We use the IEEE 39-bus system [39], represented graphically in Fig. 2, as a benchmark. Nodes denote buses, and edges represent transmission lines and transformers [3]. The parameters $\boldsymbol{\theta}$ to be inferred are the system states (voltage magnitude and angle), and $\bm{f}$ represents PMU measurements. Full parameter identifiability implies that the target QoIs are the state variables themselves [$\bm{g}(\boldsymbol{\theta};\mathbf{y})=\boldsymbol{\theta}$] and the target FIM is non-singular. We enforce this condition by setting $\mathcal{J}=\lambda I$ for some small $\lambda>0$. The information-matching condition implies that $\mathcal{I}$ is also non-singular, and all states are observable. The optimal PMU placements for this problem are represented as the orange highlighted buses in Fig. 2. Our results on this initial benchmark test align with previous studies [4, 5, 3]. Even without preassigning PMU locations, this method naturally selects the same buses identified elsewhere [6, 7].
@@
Next, we consider an optimal sensor placement problem in passive acoustic source localization in the ocean. The objective is to determine the optimal sensor (sound receiver) locations to infer the location of sound sources in a shallow ocean. This problem is difficult because sound propagation depends in complicated ways on unknown properties of the ocean envrionment, including the water temperature and the sediments in the seabed. Traditionally, an estimate of the environmental parameters of the ocean are needed to localize the source, as in matched-field processing [43]. In constrast, information-matching allows us to learn only those combinations of acoustic parameters that are necessary to infer the source location.
@@
To localize the two sources using Eq. (4), the target FIM $\mathcal{J}$ is a diagonal matrix, where diagonal elements that correspond to source locations are set to their inverse target precision, and all other elements are zero. The results indicate the the optimal receiver locations for this problem as shown as the large dots in Fig. 3, which consist of only 5% of the candidate locations.
@@
Algorithm 1 outlines the iterative AL process based on information-matching. We begin by preparing a pool of candidate inputs $\mathbf{X}=\{\mathbf{x}_{m}\}_{m=1}^{M}$ and initializing the parameters to the a priori best estimate $\boldsymbol{\theta}_{0}$. The procedure starts by evaluating $\mathcal{J}$ and $\mathcal{I}_{m}$ at $\boldsymbol{\theta}_{0}$ for all $m$ inputs and solving Eq.(4) for the weights $\mathbf{w}$. For each datum, we update the optimal weight by comparing the new optimal value with the current one, retaining the larger one. This step ensures the amount of information in the training data in subsequent iterations is non-decreasing. Convergence occurs when the change in optimal weights between subsequent iterations is below some chosen threshold. If not converged, we generate labels for data with nonzero weights and update the parameters by minimizing Eq. (1), then iterate.
@@
2:      $\mathbf{X}\leftarrow\{\mathbf{x}_{m}\}_{m=1}^{M}$triangleright Candidate input data
$\mathbf{P}\leftarrow\text{empty}(M)$triangleright To store labels
$\mathbf{w}^{\text{opt}}\leftarrow\text{zeros}(M)$triangleright To store optimal weights
$\boldsymbol{\theta}_{0}\leftarrow\boldsymbol{\theta}$triangleright Initial parameters
@@
16:         $\boldsymbol{\theta}_{0}\leftarrow\operatorname*{arg\,min}_{\boldsymbol{\theta}}\ell(\boldsymbol{\theta})$triangleright Update parameters
@@
In summary, we have introduced an information-matching method to identify a minimal set of informative data points to meet target precision requirements. This approach involves solving the convex optimization problem in Eq. (4) to align the information available in the training data with that necessary for precise predictions. Unlike other FIM-based OED criteria, this method selectively constrains only the parameters relevant for the downstream QoIs, making it robust to numerical noise in sloppy models with unidentifiable parameters and small FIM eigenvalues. The approach not only improves data efficiency, but also enhances model interpretability.
@@ Acknowledgements.
YK, MKT, IN, EBT, and AMS acknowledge partial support through NSF under Grants No. 1834251, 1834332, and 2223986. YK, IN, EBT, VVB, VL, and MKT acknowledge funding support from the Laboratory Directed Research and Development program (project code 23-SI-006) and a special computational time allocation on the Lassen supercomputer from the Computational Grand Challenge program at Lawrence Livermore National Laboratory. Portions of this work were performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344. TBN and MKT acknowledge that this work related to Department of the Navy Award N00014-24-12566. Any opinions, findings, conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect views of the Office of Naval Research. Some of the calculations were conducted using computational facilities provided by the Brigham Young University Office of Research Computing. Details about the models and additional results are provided in the SM. The data and code to reproduce the results of this study are available on Figshare [52] and GitHub [53].

### v3

@@ Abstract
Many OED criteria utilize the Fisher information matrix (FIM), whose inverse establishes a lower bound on parameter covariance, known as the Cramér-Rao bound [22, 23, 24]. Common approaches optimize parameter precision through the FIM, for example by minimizing its trace (A-optimality) [25, 26, 27], maximizing its determinant (D-optimality) [28, 29, 16, 27], or maximizing the smallest eigenvalues (E-optimality) [30, 29, 31].
However, many applications of predictive models do not require precise parameter estimates per se, but accurate predictions for key quantities of interest (QoIs) [18, 32, 33, 34, 35]. This distinction is well illustrated by sloppy models, where many parameter combinations are practically unidentifiable, yet still yield precise predictions [36, 37, 38]. Such models typically have only a few stiff, identifiable directions, while the rest are sloppy and poorly constrained. Furthermore, the parameters that are identifiable from the training data may not align with those most relevant for predicting QoIs. Precise predictions are still possible when these QoI-relevant directions fall within the identifiable subspace; however, if they do not, the resulting uncertainty may be large or even divergent, regardless of how well some parameters are constrained [39].
Furthermore, sloppy models exhibit a characteristic information spectrum, where FIM eigenvalues are nearly uniformly spaced on a log scale over many orders of magnitude. Many eigenvalues are smaller than the model’s evaluation precision, rendering the aforementioned OED criteria sensitive to numerical noise.
Motivated by these considerations, we develop an information-matching method for OED that prioritizes the precision of predictions for the target QoIs. This approach uses a model parameterized by $\boldsymbol{\theta}$ in two key scenarios: training, $\bm{f}$, and prediction, $\bm{g}$ (see Fig. 1). Given a dataset of $M$ independent inputs $\{\mathbf{x}_{m}\}_{m=1}^{M}$ and their corresponding ground truth labels $\{\mathbf{p}_{m}\}_{m=1}^{M}$, we first use $\bm{f}(\boldsymbol{\theta};\mathbf{x}_{m})$ to train the parameters against the data $\{\mathbf{x}_{m},\mathbf{p}_{m}\}$. Then, we use the trained parameters to predict the target QoIs $\mathbf{q}$ corresponding to the input $\mathbf{y}$ through the mapping $\bm{g}(\boldsymbol{\theta};\mathbf{y})$. In this scenario, $\mathbf{y}$ may act as an input, similar to $\mathbf{x}_{m}$, as a control parameter for the QoIs, or as a discrete index to distinguish between different QoI values. The information-matching method leverages both scenarios to identify the minimal subset of training data that contains the information necessary to precisely constrain the parameters relevant to target QoIs. Our strategy is: given a target precision for target QoIs, align the FIM for the training data with that of the QoIs, ensuring the training data carry the information needed to constrain the predictions precisely. Thus, only the parameter combinations that need to be identified are trained, bypassing numerical stability issues in cases where the FIMs are ill-conditioned.
@@
where the weight $w_{m}$ is the inverse variance of the label $\mathbf{p}_{m}$, i.e., $w_{m}=1/\sigma_{m}^{2}$. As shown in the supplementary material, the FIM for the training scenario is given by
@@
Let $\bm{g}(\boldsymbol{\theta};\mathbf{y})$ denote a mapping from the model parameters $\boldsymbol{\theta}$ to the target QoIs for input $\mathbf{y}$ that is analytic at $\boldsymbol{\theta}_{0}=\left<\boldsymbol{\theta}\right>_{\boldsymbol{\theta}}$, where $\left<\cdot\right>_{\boldsymbol{\theta}}$ denotes an expectation value over the distribution of parameters. Consider parameters of the form $\boldsymbol{\theta}=\boldsymbol{\theta}_{0}+\epsilon\delta\boldsymbol{\theta}$. If the constraints in Eq. (4) are satisfied, then
@@
Theorem 1 states that the uncertainties of the QoIs propagated from the optimal training data $\mathrm{Cov}(\bm{g})$] are within the predefined target uncertainties ($\mathbf{\Sigma}$), up to third order in $\epsilon$. The information-matching method is unique in that it simultaneously aims to minimize data usage while ensuring adequate information for precise predictions. In contrast, the previously mentioned OED criteria (A-, D-, and E-optimality) only prioritize reducing some measure of parameter variance. By minimizing the number of data points, the information-matching approach not only enhances efficiency but also improves model interpretability by focusing the analysis on only the most critical training data.
@@
We use the IEEE 39-bus system [40], represented graphically in Fig. 2, as a benchmark. Nodes denote buses, and edges represent transmission lines and transformers [3]. The parameters $\boldsymbol{\theta}$ to be inferred are the system states (voltage magnitude and angle), and $\bm{f}$ represents PMU measurements. Full parameter identifiability implies that the target QoIs are the state variables themselves $\bm{g}(\boldsymbol{\theta};\mathbf{y})=\boldsymbol{\theta}$] and the target FIM is non-singular. We enforce this condition by setting $\mathcal{J}=\lambda I$ for some small $\lambda>0$. The information-matching condition implies that $\mathcal{I}$ is also non-singular, and all states are observable. The optimal PMU placements for this problem are represented as the orange highlighted buses in Fig. 2. Our results on this initial benchmark test align with previous studies [4, 5, 3]. Even without preassigning PMU locations, this method naturally selects the same buses identified elsewhere [6, 7].
@@
Next, we consider an optimal sensor placement problem in passive acoustic source localization in the ocean. The objective is to determine the optimal sensor (sound receiver) locations to infer the location of sound sources in a shallow ocean. This problem is difficult because sound propagation depends in complicated ways on unknown properties of the ocean environment, including the water temperature and the sediments in the seabed. Traditionally, an estimate of the environmental parameters of the ocean is needed to localize the source, as in matched-field processing [44]. In contrast, information-matching allows us to learn only those combinations of acoustic parameters that are necessary to infer the source location.
@@
To localize the two sources using Eq. (4), the target FIM $\mathcal{J}$ is a diagonal matrix, where diagonal elements that correspond to source locations are set to their inverse target precision, and all other elements are zero. The optimal receiver locations are shown as the large dots in Fig. 3, which account for only 5% of the candidate locations.
@@
2:   $\mathbf{X}\leftarrow\{\mathbf{x}_{m}\}_{m=1}^{M}$ $\triangleright$ Candidate input data
$\mathbf{P}\leftarrow\text{empty}(M)$ $\triangleright$ To store labels
$\mathbf{w}^{\text{opt}}\leftarrow\text{zeros}(M)$ $\triangleright$ To store optimal weights
$\boldsymbol{\theta}_{0}\leftarrow\boldsymbol{\theta}$ $\triangleright$ Initial parameters
@@
16:   $\boldsymbol{\theta}_{0}\leftarrow\operatorname*{arg\,min}_{\boldsymbol{\theta}}\ell(\boldsymbol{\theta})$ $\triangleright$ Update parameters
@@
Algorithm 1 outlines the iterative AL process based on information-matching. We begin by preparing a pool of candidate inputs $\mathbf{X}=\{\mathbf{x}_{m}\}_{m=1}^{M}$ and initializing the parameters to the a priori best estimate $\boldsymbol{\theta}_{0}$. The procedure starts by evaluating $\mathcal{J}$ and $\mathcal{I}_{m}$ at $\boldsymbol{\theta}_{0}$ for all $m$ inputs and solving Eq.(4) for the weights $\mathbf{w}$. For each datum, we update the optimal weight by comparing the new optimal value with the current one, retaining the larger one. This step ensures that the amount of information in the training data in subsequent iterations is non-decreasing. Convergence occurs when the change in optimal weights between subsequent iterations is below some chosen threshold. If not converged, we generate labels for data with nonzero weights and update the parameters by minimizing Eq. (1), then iterate.
@@
In summary, we have introduced an information-matching method to identify a minimal set of informative data points to meet target precision requirements. This approach involves solving the convex optimization problem in Eq. (4) to align the information available in the training data with that necessary for precise predictions. Unlike other FIM-based OED criteria, this method selectively constrains only the parameters relevant for the downstream QoIs, making it robust to numerical noise in sloppy models with unidentifiable parameters and small FIM eigenvalues. The approach not only improves data efficiency but also enhances model interpretability.
@@
Supplementary Material
The supplementary material provides the mathematical formulation of the least-squares problem and the associated uncertainty quantification framework. It also includes a proof of the information-matching theorem, detailed descriptions of the models used in the example test cases, and additional results for those models.
@@ Acknowledgements.
YK, MKT, IN, EBT, and AMS acknowledge partial support through NSF under Grants No. 1834251, 1834332, and 2223986. YK, IN, EBT, VVB, VL, and MKT acknowledge funding support from the Laboratory Directed Research and Development program (project code 23-SI-006) and a special computational time allocation on the Lassen supercomputer from the Computational Grand Challenge program at Lawrence Livermore National Laboratory. Portions of this work were performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344. TBN and MKT acknowledge that this work is related to the Department of the Navy Award N00014-24-12566. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the Office of Naval Research. Some of the calculations were conducted using computational facilities provided by the Brigham Young University Office of Research Computing.
Data Availability Statement
The data that support the findings of this study are openly available in Figshare at http://doi.org/10.6084/m9.figshare.27274029.v1, reference number [52]. The corresponding code is available in Github, reference number [53].
AUTHOR DECLARATION
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
Yonatan Kurniawan: Conceptualization (equal); Formal Analysis (equal); Investigation (lead); Methodology (equal); Software (lead); Writing—original draft (lead). Tracianne Neilsen: Data curation (equal); Writing—original draft (equal). Benjamin Francis: Data curation (equal); Writing—original draft (equal). Alex Stankovic: Writing—review & editing (equal). Mingjian Wen: Data curation (equal); Writing—review & editing (equal). Ilia Nikiforov: Writing—review & editing (supporting). Ellad Tadmor: Writing—review & editing (equal). Vasily Bulatov: Writing—review & editing (equal). Vincenzo Lordi: Writing—review & editing (equal). Mark Transtrum: Conceptualization (equal); Formal analysis (equal); Methodology (equal); Writing—original draft (equal).